The volume of a certain pipe varies directly with its length. The table below shows the volume of the pipe at two lengths. Find the volume of the pipe at the third length in the table.
Volume 30 gallons 45 gallons ?
Length 20 feet 30 feet 40 feet

Answer by edjones(8007)

x/40=45/30 Proportion.
.
x=45*40/30=60 gallons
.
Ed